I set the MC back to its good alignment (June 21st) using this procedure. The trend of the OSEM values over the last 40 days and 40 nights is attached.
Then I aligned the periscope to that beam. This took some serious periscope knob action. Without WFS, the transmission went to 2.7 V and the reflection down to 0.6V.
Then I re-aligned the MC_REFL path as usual. The beam was far enough off that I had to also re-align onto the MC LSC PD as well as the MC REFL camera (~2 beam radii).
Beams are now close to their historical positions on Faraday and MC2. I then restored the PZT sliders to their April snapshot and the X-arm locked.
Steve - please recenter the iris which is on the periscope. It has been way off for a long time.
So it looks OK now. The main point here is that we can trust the MC OSEMs.
Afterwards I rebooted c1susvme1 and c1susvme2 because they were skewed.
It is really surprising that we now have again the data from the MC OSEMs since up to two days ago the record looked corrupted (see the attachments in my entry 1774).
The reason I ended up severely misaligning the the MC is exactly that there wasn't anymore a reference position that I could go back to and I had to use the camera looking a the Faraday.